OpenStudy (anonymous):

apply the Rational Root Theorem to f(x)=2x^4+15x^3+26x^2-5x-6

OpenStudy (anonymous):

\[2 x^4+15 x^3+26 x^2-5 x-6=(x+3) (2 x-1) \left(x^2+5 x+2\right) \]

OpenStudy (anonymous):

how can i use that to find possible rational roots?

OpenStudy (anonymous):

Set each factor on the RHS and solve for x.

OpenStudy (anonymous):

RHS?

OpenStudy (anonymous):

RHS means is an abbreviation for "right hand side" LHS is short for "Left hand side"

OpenStudy (anonymous):

If x+3 is zero, then x = -3. Evaluate the problem expression with x = -3 and the result should be zero.

OpenStudy (anonymous):

so the answer is -3, 1/2, and x = -0.2*x^2-0.4?

OpenStudy (anonymous):

The solution to \[x^2+5 x+2=0\]is\[\left\{x=\frac{1}{2} \left(-5-\sqrt{17}\right),x=\frac{1}{2} \left(-5+\sqrt{17}\right)\right\} \]You can use the binomial theorem to solve it.

OpenStudy (anonymous):

The first two solutions, -3 and 1/2 are correct.